### v3

Abstract
This study considers the method to derive a ranking of alternatives by aggregating the rankings submitted by several individuals who may not evaluate all of them. The collection of subsets of alternatives that individuals (can) evaluate is referred to as an evaluability profile. For a given evaluability profile, we define an aggregating ranking function whose inputs are the rankings of individuals on the alternatives they evaluate. We investigate the properties of aggregating ranking functions, which are modifications of those introduced in previous studies. Whether an aggregating ranking function satisfying a combination of properties exists depends on the evaluability profile. Thus, we identify the necessary and sufficient conditions on evaluability profiles to ensure the existence of the functions satisfying four different combinations of properties. Furthermore, to examine how frequently possible or impossible evaluability profiles occur, we derive the proportion of each type in specific cases.
JEL classification codes: D71; D72
Keywords: Aggregating ranking function; Social welfare function; Evaluability profile; Pareto principle; Arrow’s impossibility result; Peer rating
1 Introduction
This study examines situations in which the ranking of finite alternatives is required, involving the rankings of these alternatives by a number of individuals. Examples illustrating this situation include a faculty considering several candidates for new members, or a company evaluating various project proposals. In such organizations, decisions are often influenced by the evaluation rankings submitted by their members. Thus, the planner asks individuals to submit their rankings of alternatives, but the rankings submitted by some individuals may be incomplete. In a real-world example, the evaluation of a specific competitive research grant is conducted through peer review. During the review process, evaluations of researchers with whom there is a conflict of interest are designed to be avoided through self-disclosure.
The reasons why some individuals do not or should not submit complete ranking are such as those as follows. First, they may lack sufficient information or knowledge to evaluate these alternatives. For example, evaluating potential faculty members may require expertise in their specific field. Second, an individual is known to be biased in favor of their family members, colleagues, themselves, their own ideas, and the projects they were previously involved in. [^1] In psychology, such a bias is called “ingroup bias.” Similar biases are also frequently observed in the behavioral economics literature. For instance, several studies are introduced in Ariely’s (2010) third and fourth chapters. Moreover, there are several voting models that consider the biases of voters. For example, Amorós et al. (2002), Amorós (2009, 2011, 2019) and Adachi (2014) consider a jury problem where some juries have some biases. As a result, they may not be well-suited to provide objective evaluations of these alternatives. Third, planners may encounter difficulties in obtaining complete rankings from all individuals. For example, when new alternatives are added after some individuals have already submitted their rankings, there may be no opportunity to gather their rankings for the newly added ones. Consequently, the evaluations by these individuals may not be incorporated into the aggregated ranking of the alternatives.
To address such situations, we revisit a model originally introduced by Ohbo et al. (2005) and also discussed by Ando et al. (2007). In the model, each individual is assigned a subset of alternatives to evaluate and is asked to submit a weak order (ranking) of them. In this study, we refer to the collection of these subsets as an evaluability profile. [^2] Several previous studies, including Quesada (2005) and Cengelci and Sanver (2007) examine models in which voters have the option to abstain from voting. In contrast, this study assumes that the evaluability profile is independent of individuals’ choices and is exogenously given. Given an evaluability profile, we define an aggregating (incomplete) ranking function that takes the rankings submitted by all individuals as inputs. This function is equivalent to the social welfare function à la Arrow (1963) when the ranking of each individual is complete. Moreover, in the peer rating model introduced by Ando et al. (2003) and Ng and Sun (2003), individuals evaluate each other, but no individual evaluates themselves. [^3] The prize award problem, as discussed by Holzman and Moulin (2013), Tamura and Ohseto (2014), Mackenzie (2015, 2019), Tamura (2016) and Edelman and Por (2021) is related to the peer rating problem. Our model is also a generalization of the peer rating model.
We examine six properties of aggregating ranking functions: transitive valuedness, the Pareto criterion, the weak Pareto criterion, independence of irrelevant alternatives, non-dictatorship, and non-consistency which are modifications of those of Arrow (1963) and Hansson (1973).
First, we derive the necessary and sufficient condition on the evaluability profile to ensure the existence of an aggregating ranking function that satisfies transitive valuedness and the Pareto criterion. Second, we additionally require independence of irrelevant alternatives but show that the necessary and sufficient condition is unchanged. Third, we derive the necessary and sufficient condition ensuring the existence of an aggregating ranking function that satisfies non-dictatorship in addition to the properties of the second result. We need an additional condition for the existence of an aggregating ranking function that additionally satisfies non-dictatorship. Fourth, we revisit the properties in the second result but the weak Pareto criterion and non-consistency are considered instead of the Pareto criterion, because the Pareto criterion implies them. However, the necessary and sufficient condition is the same as that of the second result as well as the first result.
The first result introduced above is a generalization of that of Ando et al. (2003) which specifically focus on the peer rating model. Moreover, the fourth result introduced above is also closely related to one of the results of Ando et al. (2003). The proofs of these results are inspired by their proof techniques. Furthermore, the third result is a generalization of that of Ohbo et al. (2005) that consider the case involving at least one individual who evaluates all alternatives. While the results of Ando et al. (2003) and Ohbo et al. (2005) show impossibility only, we introduce specific ARFs to have the conditions that are sufficient for the existence of ARFs.
By the second and third results, we can distinguish evaluability profiles into three types. The first type is characterized by the absence of ARFs that satisfy transitive valuedness, the Pareto criterion, and independence of irrelevant alternatives. The second type is such that an ARF satisfying the three must be a dictatorship rule. The third type is one for which there is an ARF that satisfies non-dictatorship in addition to the three properties. To examine how frequently each type occurs, we derive the proportion of each type in specific cases.
There are three recent related studies to ours. First, Fujiwara-Greve et al. (2023) independently discuss a model similar to ours. In their model, described in our terminology, an evaluability profile is also an input of the aggregating ranking function. They consider several combinations of properties on the function, which differ from ours, and discuss which combinations are possible or impossible for the existence of a function satisfying them. On the other hand, in our model, for a given evaluability profile, the aggregating ranking function is defined, and we consider the conditions on the evaluability profiles. [^4] Since in the combinations of axioms in this study whether impossible or possible is dependent on the evaluability profile, these combinations are “impossible” in the sense of Fujiwara-Greve et al. (2023). Moreover, Fujiwara-Greve et al. (2023) restrict their attention to a class of evaluability profiles. In this study, we also clarify the results when we focus on the class that they consider. Second, Barberà and Bossert (2023) introduce a general model including the case where the rankings submitted by individuals may be incomplete. In their approach, they aggregate the opinions of individuals without identifying who submitted them. On the other hand, since we require non-dictatorship in our analysis, we need to identify whose opinions they are. Villar (2023) also introduces a similar general model that includes the case of incomplete preferences and introduces aggregating functions that provide a quantitative assessment to each alternative and satisfy a combination of properties. On the other hand, our aggregating functions provide just a weak order of alternatives, and thus their properties also differ from ours.
2 Model and technical results
2.1 Problems
We consider a model that is originally introduced by Ohbo et al. (2005). Let $A$ be the finite set of alternatives. Let $V$ be the finite set of individuals (voters) who submit a ranking of alternatives. We assume $\left|A\right|\geq 3$ and $\left|V\right|\geq 3$.
We consider the situation where some individuals may not evaluate some of the alternatives. Let $A_{v}\subseteq A$ be an evaluable set of individual $v$ representing the set of alternatives that $v$ (can) evaluate and $\mathbf{A}=\left(A_{v}\right)_{v\in V}\subseteq A^{\left|V\right|}$ be an evaluability profile. We assume that for any $v\in V$, at least two alternatives are in $A_{v}$. For $a\in A,$ let
$V\left(a\right)=\left\{v\in V\text{ }\left|\text{ }a\in A_{v}\right.\right\}\text{,}$
which is the set of individuals who (can) evaluate $a$. If $A_{v}=A$ (or equivalently if $v\in V\left(a\right)$ for all $a\in A$), then individual $v$ is said to be complete.
An evaluability profile $\mathbf{A}=\left(A_{v}\right)_{v\in V}$ is said to be larger (resp. smaller) than another $\mathbf{A}^{\prime}=\left(A_{v}^{\prime}\right)_{v\in V}$ if $A_{v}^{\prime}\subseteq A_{v}$ (resp. $A_{v}\subseteq A_{v}^{\prime}$) for all $v\in V$ and $A_{v}^{\prime}\subsetneq A_{v}$ (resp. $A_{v}\subsetneq A_{v}^{\prime}$) for some $v\in V$.
We say that a problem is an ordered tuplet of the sets of individuals and alternatives, and an evaluability profile $(V,A,\mathbf{A})$, which are the primitives of this study. Hereafter, we fix a problem $(V,A,\mathbf{A})$.
If $A_{v}=A$ for all $v\in V$, then the problem is equivalent to that of the Arrow’s (1963) classical problem. Moreover, suppose that $A=V;$ that is, the sets of individuals and alternatives are equivalent, and $\mathbf{A}$ is such that $A\setminus A_{v}=\left\{v\right\}$ for all $v\in V$. Then, the problem is equivalent to the peer rating problem, introduced by Ando et al. (2003) and Ng and Sun (2003); that is, which considers the situation where the individuals evaluate each other but each individual does not evaluate themselves.
Fujiwara-Greve et al. (2023) call an evaluability profile $\mathbf{A}$ is non-trivial if for any pair of alternatives $a,b\in A$, there is at least one $v\in V$ such that $a,b\in A_{v}$; that is, there is no pair of alternatives that are evaluated by nobody. They focus only on the non-trivial evaluability profiles.
2.2 Binary relations
For $A^{\prime}\subseteq A$, $R\subseteq A^{\prime}\times A^{\prime}$ be a binary relation on $A^{\prime}$. If $\left(a,b\right)\in R$, then we write $aRb$. Let $P\left(R\right)\,$and $I\left(R\right)\,$be the asymmetric part and the symmetric part of a binary relation on $A^{\prime}$ denoted by $R;$ that is,
$P\left(R\right)$ $=$ $\left\{\left.\left(a,b\right)\in A^{\prime}\times A^{\prime}\text{ }\right|\text{ }aRb\text{ and }\lnot\left(bRa\right)\right\},$
$I\left(R\right)$ $=$ $\left\{\left.\left(a,b\right)\in A^{\prime}\times A^{\prime}\text{ }\right|\text{ }aRb\text{ and }bRa\right\}.$
A binary relation $R$ on $A^{\prime}\subseteq A$ is
reflexive if for all $a\in A^{\prime}$, $aRa$,
complete if for all $a,b\in A^{\prime},$ $aRb$ or $bRa$ (or both),
transitive if for all $a,b,c\in A^{\prime},$ [$aRb$ and $bRc$] implies $aRc$,
anti-symmetric if for all $a,b\in A^{\prime},$ $aI\left(R\right)b$ implies $a=b$,
asymmetric if for all $a,b\in A^{\prime},$ $aRb$ implies $\lnot\left(bRa\right)$,
acyclic if, for all $M\in\mathbb{N}\setminus\{1\}$ and all $a_{1},a_{2},\cdots,a_{M}\in A^{\prime}$, $a_{m+1}Ra_{m}$ for all $m\in\left\{1,\cdots,M-1\right\}$ imply $\lnot\left(a_{1}Ra_{M}\right)$.
See, for example, Duggan (1999) for a detailed explanations on them. Note that if $R$ (on $A^{\prime}$) is acyclic, then it is asymmetric.
A binary relation $R$ on $A^{\prime}$ is said to be a weak order on $A^{\prime}$ if it is reflexive, complete and transitive. A binary relation $R$ on $A^{\prime}$ is said to be a linear order on $A^{\prime}$ if it is reflexive, complete, transitive and anti-symmetric. If $R$ is a linear order on $A^{\prime}$, then we write
$R:a_{1},a_{2},\cdots,a_{\left|A^{\prime}\right|}$
representing $a_{x}Ra_{x+1}$ for all $x=1,\cdots,\left|A^{\prime}\right|-1$.
A binary relation $R^{\prime}$ on $A^{\prime}\subseteq A$ is a linear order extension of an asymmetric binary relation $R$ on $A^{\prime}$ if $R^{\prime}$ is a linear order on $A^{\prime}$ and $R\subseteq R^{\prime}$. Since $R$ is asymmetric, a linear order extension of $R$ is well-defined. Let $\mathcal{E}\left(R\right)$ be the set of all linear order extensions of $R$ on $A^{\prime}$. The following fact is also known.
Remark 1
If $R$ on $A^{\prime}\subseteq A$ is acyclic, then $\mathcal{E}\left(R\right)\neq\emptyset$. [^5] If $A$ is an inifinite set, then acyclicity is not sufficient for the existence of linear order extenstions. In this case, Suzumura-consistency is a necessary and sufficient condition for the existence of linear order extenstions. See, Theorem 2.6 of Bossert and Suzumura (2010), which is a generalization of Szplrajn’s theorem. However, if $A$ is a finite set, then acyclicity is sufficient.
Note that since an acyclic binary relation is asymmetric, If $R$ on $A^{\prime}\subseteq A$ is acyclic, then a linear order extension $R$ is well-defined.
A linear order extension of an acyclic $R$ on $A^{\prime}$ is constructed as follows. First, choose one $a_{1}\in A^{\prime}$ such that $\left(a,a_{1}\right)\notin R$ for all $a\in A^{\prime}$. A result introduced by Bossert and Suzumura (2010, Theorem 2.4) ensures the existence of such $a_{1}$ with acyclic $R$. Similarly, for $t=2,\cdots,\left|A^{\prime}\right|$, we can let $a_{t}\in A^{\prime}\setminus\left\{a_{1},\cdots,a_{t-1}\right\}$ such that $\left(a,a_{t}\right)\notin R$ for all $a\in A^{\prime}\setminus\left\{a_{1},\cdots,a_{t-1}\right\}$. Then, $R:a_{1},a_{2},\cdots,a_{\left|A^{\prime}\right|}$ is a linear order extension of an acyclic $R$ on $A^{\prime}$. See also Kitahara and Okumura (2023) for this construction.
Finally, for a binary relation on $A^{\prime}\subseteq A$ denoted by $R$ and $A^{\prime\prime}\subseteq A^{\prime}$, let
$\left.R\right|_{A^{\prime\prime}}=R\cap\left(A^{\prime\prime}\times A^{\prime\prime}\right),$
which is a binary relation on $A^{\prime\prime}$.
2.3 Aggregating ranking functions and properties
Let $\mathcal{W}_{v}$ be the set of all possible weak orders on $A_{v}$. Let $R_{v}\in\mathcal{W}_{v}$ represent a ranking on $A_{v}$ submitted by $v$ [^6] A voter $i$ may have an evaluation order $R_{v}^{\prime}$ on $A^{\prime}\supsetneq A_{v},$ but he/she reveals a preference order on $A_{v}$. In that case, $R_{v}$ $=\left.R_{v}^{\prime}\right|_{A_{v}}$. Moreover, let
$\mathcal{W\equiv}\prod\limits_{v\in V}\mathcal{W}_{v}\text{ and }\mathbf{R}\left(\mathcal{\equiv}\left(R_{v}\right)_{v\in V}\right),\mathbf{R}^{\prime}\left(\mathcal{\equiv}\left(R_{v}^{\prime}\right)_{v\in V}\right)\in\mathcal{W}.$
Let $f:\mathcal{W}\rightarrow\mathcal{B}$ be an aggregating ranking function (hereafter ARF), where $\mathcal{B}$ is the set of all reflexive and complete binary relations on $A$.
We introduce the properties of $f$ that are modifications of those introduced by Arrow (1963). First, an ARF $f$ satisfies transitive valuedness (hereafter TV) if $f\left(\mathbf{R}\right)$ is transitive for all $\mathbf{R}\in\mathcal{W}$.
Second, we say that an ARF $f$ satisfies the Pareto criterion (hereafter PC), if for all $\mathbf{R}\in\mathcal{W}$, and all $a,b\in A,$ if $aP\left(R_{v}\right)b$ for all $v\in V\left(a\right)\cap V\left(b\right)$, then $aP\left(f\left(\mathbf{R}\right)\right)b$.
Third, we say that $f$ satisfies the weak Pareto criterion (hereafter wPC), if for all $\mathbf{R}\in\mathcal{W}$, and all $a,b\in A,$ if $aP\left(R_{v}\right)b$ for all $v\in V\left(a\right)\cap V\left(b\right)$, then $af\left(\mathbf{R}\right)b$ (but $bf\left(\mathbf{R}\right)a$ is allowed to hold). Ando et al. (2003) and Ohseto (2007) consider this in the peer rating model.
Fourth, we say that $f$ satisfies independence of irrelevant alternatives (hereafter IIA), if for all $a,b\in A$ and all $\mathbf{R,R}^{\prime}\in\mathcal{W},$ $\left.R_{v}\right|_{\{a,b\}}=\left.R_{v}^{\prime}\right|_{\{a,b\}}$ for all $v\in V\left(a\right)\cap V\left(b\right)$ implies $\left.f\left(\mathbf{R}\right)\right|_{\{a,b\}}=\left.f\left(\mathbf{R}^{\prime}\right)\right|_{\{a,b\}}$. Note that for $a,b\in A$ such that $V\left(a\right)\cap V\left(b\right)=\emptyset$, $\left.f\left(\mathbf{R}\right)\right|_{\{a,b\}}$ is dependent on irrelevant alternatives, because in this case, there is no individual who evaluate both $a$ and $b$.
Fifth, we define a “dictator” of this model. First, individual $v$ is said to be a quasi-dictator of $f$ if for all $\mathbf{R}\in\mathcal{W}$ and all $a,b\in A,$ $aP\left(R_{v}\right)b$ implies $aP\left(f\left(\mathbf{R}\right)\right)b$. Second, individual $v$ is said to be a dictator of $f$ if $v$ is a quasi-dictator of $f$ and complete. If $v$ is not complete, then the positions of the alternatives in $A\setminus A_{v}$ are determined without any opinion of $v$. Thus, in that case, $v$ is not called a dictator of $f$, even if the rankings revealed by $v$ is always reflected in $f$. That is, in this study, a quasi-dictator is a dictator only if they evaluates all alternatives. We say that $f$ satisfies non-dictatorship (hereafter ND) if there is no dictator of $f$. Note that if there is no complete individual , then any ARF satisfies ND, because nobody can be a dictator for that profile in our definition.
Finally, we say that an ARF $f$ satisfies non-constancy (hereafter NC), if there is no pair of alternatives $\left(a,b\right)$ with $a\neq b$ and $V\left(a\right)\cap V\left(b\right)\neq\emptyset$ such that $aP\left(f\left(\mathbf{R}\right)\right)b$ for all $\mathbf{R}\in\mathcal{W}$ (or $bP\left(f\left(\mathbf{R}\right)\right)a$ for all $\mathbf{R}\in\mathcal{W}$) or $aI\left(f\left(\mathbf{R}\right)\right)b$ for all $\mathbf{R}\in\mathcal{W}$. This property is a modification of that introduced by Hansson (1973). Ando et al. (2003) also consider this property in the peer rating model.
Remark 2
If an ARF $f$ satisfies PC, then it also satisfies wPC and NC.
By this remark, when PC is required, we do not explicitly require wPC and NC. However, wPC does not imply NC, and vice versa.
Note that except for wPC, ND, and NC, the definitions of the properties are the same as those introduced by Ohbo et al. (2005). They do not consider wPC or NC and do not explicitly define ND. However, since they show the existence of a dictator, who is a quasi-dictator and a complete individual, they (implicitly) focus on ND in our definition.
2.4 Graphs
Since we use some graph theory descriptions to introduce our results, we first present some notations of graph theory. Throughout of this paper, we let the set of alternatives $A$ be the set of nodes. An (undirected) graph is a set of unordered pairs of nodes. If $\left\{a,b\right\}$ is an element of a graph $G$, then we simply write $ab\in G$ or equivalently $ba\in G$. If $ab\in G$ for all $a,b\in A$ such that $a\neq b$, then $G$ is said to be the complete graph.
A subset of graph $G$ denoted by $P\subseteq G$ is said to be an (undirected) path of $G$ if
$P=\left\{a_{1}a_{2},a_{2}a_{3},\cdots,a_{M-1}a_{M}\right\},$
where $a_{1},a_{2},\cdots,a_{M}$ are distinct. A subset of graph $G$ denoted by $C\subseteq G$ is said to be an (undirected) cycle of $G$ if
$C=\left\{a_{1}a_{2},a_{2}a_{3},\cdots,a_{M-1}a_{M},a_{M}a_{1}\right\},$ (1)
where $a_{1},a_{2},\cdots,a_{M}$ are distinct and $M\geq 3$, which rules out that $C=\left\{a_{1}a_{2},a_{2}a_{1}\right\}$ is a cycle. For notational convenience, when we use a cycle denoted by (1), $a_{M+1}=a_{1}$. If all nodes are included in a cycle; that is, if $\left\{a_{1},a_{2},\cdots,a_{M}\right\}=A$, then $C$ is said to be a Hamiltonian cycle of $G$. Trivially, the complete graph has Hamiltonian cycles.
We consider two cycles $C$ and $C^{\prime}$ of $G,$ where the nodes included in $C$ and $C^{\prime}$ are represented by $A$ and $A^{\prime}$, respectively. We say that $C$ is larger than $C^{\prime}$ if $A^{\prime}\subsetneq A$. A cycle $C$ of $G$ is said to be a maximal cycle of $G$ if there is no cycle $C^{\prime}$ of $G$ that is larger than $C$.
We construct graphs from a given evaluability profile $\mathbf{A}$. For $v\in V$, let
$G_{v}=\left\{ab\text{ }\left|\text{ }a,b\in A_{v}\text{ and }a\neq b\right.\right\};$
that is, if individual $v$ evaluates both $a$ and $b$, then $ab\in G_{v}$. Moreover,
$G=\left\{ab\text{ }\left|\text{ }a,b\in A_{v}\text{ for some }v\in V\text{ and}a\neq b\right.\right\}.$
Note that $G_{v}\subseteq G$ for all $v\in V$; that is, if there is at least one individual who evaluates both $a$ and $b$, then $ab\in G$. If there is a complete individual, then $G$ is the complete graph. On the other hand, in the peer rating model, although there is no complete individual, $G$ is the complete graph. Moreover, an evaluability profile $\mathbf{A}$ is non-trivial if and only if $G$ is the complete graph
3 Results
To introduce our results, we use the following two conditions.
Condition 1
If there are some cycles of $G$, then for each cycle of $G$ denoted by $C$, there is $v\in V$ such that $C\subseteq G_{v}$.
Condition 1 requires that for each cycle of $G$, there is at least one individual who evaluates all alternatives that are included in the cycle. Note that if there is a complete individual, then Condition 1 is satisfied. On the other hand, in the peer rating model, Condition 1 is not satisfied, because there is a Hamiltonian cycle but no complete individual.
Condition 2
There is no Hamiltonian cycle of $G$.
Condition 2 requires that $\mathbf{A}$ is sufficiently small as not to have any Hamiltonian cycle.
We have the following result on the conditions.
Lemma 1
Conditions 1 is satisfied and Condition 2 is not satisfied if and only if there is at least one complete individual.
Proof. First, we show the if-part. Suppose that there is at least one complete individual denoted by $v\in V$. Then, $G$ is the complete graph and thus has Hamiltonian cycles. Therefore, Condition 2 is not satisfied. Moreover, each cycle of $G$ denoted by $C,$ $C\subseteq G_{v}$, because $v$ is complete individual.
Second, we show the only-if-part. Suppose that Conditions 1 is satisfied and Condition 2 is not satisfied. Then, there is a Hamiltonian cycle. Since the Hamiltonian cycle includes all alternatives and Condition 1 holds, there must be at least one complete individual. Q.E.D.
We use this result to show several results introduced later.
Now, we begin deriving the condition that ensures the existence of ARFs that satisfy four different combinations of the properties introduced above. First, we consider TV and PC as properties of ARFs.
We introduce an ARF $f^{\ast}$ constructed as follows. First, let $R^{\ast}\left(\mathbf{R}\right)$ be a binary relation such that $aR^{\ast}\left(\mathbf{R}\right)b$ if and only if $V\left(a\right)\cap V\left(b\right)\neq\emptyset$ and $aP\left(R_{v}\right)b$ for all $v\in V\left(a\right)\cap V\left(b\right)$. Trivially, $R^{\ast}\left(\mathbf{R}\right)$ is asymmetric. Second, let
$f^{\ast}\left(\mathbf{R}\right)$$\in$ $\mathcal{E}\left(R^{\ast}\left(\mathbf{R}\right)\right)\text{ if }R^{\ast}\left(\mathbf{R}\right)\text{ is acyclic,}$,
$=$ arbitrary weak order if otherwise.
By Remark 1, $\mathcal{E}\left(R^{\ast}\left(\mathbf{R}\right)\right)\neq\emptyset$ if $R^{\ast}\left(\mathbf{R}\right)$ is acyclic. Thus, $f^{\ast}$ is well-defined.
Then, we have the following result.
Theorem 1
There is an ARF that satisfies TV and PC if and only if Condition 1 is satisfied.
Proof. First, we show the only-if-part. The proof of this part is inspired by the proof of Ando et al. (2003, Theorem 3.4). Suppose that Condition 1 is not satisfied; that is, there is a cycle of $G$ that is defined by (1) without $v\in V$ such that $C\subseteq G_{v}$. Moreover, let $A^{C}=\left\{a_{1},a_{2},\cdots,a_{M}\right\}\subseteq A$, which is the set of alternatives that are included in $C$.
We define a specific preference profile of individuals. First, for each $m=1,\cdots,M$, let $R^{m}$ be a linear order on $A^{C}$ that satisfies
$R^{m}:a_{m},a_{m-1},\cdots,a_{1},a_{M},a_{M-1},\cdots,a_{m+1}.$
Then, let $\mathbf{\bar{R}=}\left(\bar{R}_{v}\right)_{v\in V}$ be a cyclic profile of $C$ if for all $v\in V$, $\bar{R}_{v}$ is a linear order on $A_{v}$ satisfying
$\left.\bar{R}_{v}\right|_{A^{C}}=\left.R^{m}\right|_{A_{v}\cap A^{C}}$
for some $m=1,\cdots,M$ such that $a_{m}\notin A_{v}$; that is, $a_{m}$ is an alternative that $v$ does not evaluate. [^7] Note that Ando et al. (2003) also consider a cyclic profile only in the case where $A=V$ and $A\setminus A_{v}=\left\{v\right\}$ for all $v\in V$. This profile is well-defined for any $C\subseteq G$, because there is no $v\in V$ such that $C\subseteq G_{v}$ due to Condition 1.
[…]
Now, toward a contradiction, suppose that $f$ is TV and PC. Then, since $f$ is PC and the fact above holds, for all $m=1,\cdots,M-1,$ $a_{m+1}P\left(f\left(\mathbf{\bar{R}}\right)\right)a_{m}$ and $a_{1}P\left(f\left(\mathbf{\bar{R}}\right)\right)a_{M}$. Since $f$ is TV and $P\left(f\left(\mathbf{\bar{R}}\right)\right)$ is transitive, $a_{m+1}P\left(f\left(\mathbf{\bar{R}}\right)\right)a_{m}$ for all $m=1,\cdots,M-1$ implies $a_{M}f\left(\mathbf{\bar{R}}\right)a_{1}$ that contradicts $a_{1}P\left(f\left(\mathbf{\bar{R}}\right)\right)a_{M}$.
Second, we show the if-part. We show the following result.
Claim 1
If Condition 1 is satisfied, then $R^{\ast}\left(\mathbf{R}\right)$ is acyclic for any $\mathbf{R}\in\mathcal{W}$.
Proof. Suppose not; that is, for some $\mathbf{R}\in\mathcal{W}$, $R^{\ast}\left(\mathbf{R}\right)$ has a cycle for $A$; that is, there is some $M\in\mathbb{N}\setminus\left\{1\right\}$ and $a_{1},\cdots,a_{M}$ such that $a_{m+1}R^{\ast}\left(\mathbf{R}\right)a_{m}$ for all $m\in\left\{1,\cdots,M-1\right\}$ and $a_{1}R^{\ast}\left(\mathbf{R}\right)a_{M}$. Since $R^{\ast}$ is asymmetric, $M\geq 2$ and $a_{m+1}P\left(R^{\ast}\left(\mathbf{R}\right)\right)a_{m}$ for all $m\in\left\{1,\cdots,M\right\}$ where $a_{M+1}=a_{1}$. Then, since $a_{m+1}R^{\ast}\left(\mathbf{R}\right)a_{m}$ implies $V\left(a_{m+1}\right)\cap V\left(a_{m}\right)\neq\emptyset$, there is a cycle of $G$ denoted by (1). By Condition 1, there is $v\in V$ such that $C\subseteq G_{v}$; that is, $v\in V\left(a_{m+1}\right)\cap V\left(a_{m}\right)$ for all $m\in\left\{1,\cdots,M\right\}$. By the definition of $R^{\ast}\left(\mathbf{R}\right),$ $a_{m+1}P\left(R_{v}\right)a_{m}$ for all $m\in\left\{1,\cdots,M\right\}$, which contradicts that $R_{v}$ is a weak order. Therefore, $R^{\ast}$ is acyclic. Q.E.D.
By Remark 2 and Claim 1, if Condition 1 is satisfied, then there is a linear order extension of $R^{\ast}\left(\mathbf{R}\right)$. Therefore, in this case, $f^{\ast}$ satisfies TV and PC. Q.E.D.
Theorem 1 is a generalization of a result of Ando et al. (2003), which discuss the peer rating model and thus have an impossibility result only. This is because Condition 1 is not satisfied in the model and thus there is no ARF that satisfies TV and PC.
Second, we consider an ARF that satisfies TV, PC, and IIA. We assume that Condition 1 is satisfied, because this is a necessary condition for the existence of ARFs that satisfies TV and PC. We introduce the following ARF denoted by $f^{\ast\ast}$, which is well defined under Condition 1.
We first briefly introduce points of the ARF introduced below and denoted by $f^{\ast\ast}$. If there are some cycles of $G$, then we choose one local dictator of the alternatives included in each cycle. Condition 1 guarantees that choosing a local dictator for each cycle is possible. Moreover, as shown later, we can properly choose the local dictators to avoid any overlap.
[…]
neither $A^{x}\setminus A^{y}$ nor $A^{y}\setminus A^{x}$ is empty for all $x,y$ such that $x\neq y,$ and
if $a$ is not included in any of $C^{1},\cdots,C^{X}$ ; that is, if
$a\in A\setminus\left(A^{1}\cup\cdots\cup A^{X}\right)\left(\equiv A^{0}\right)\text{,}$
then $a$ is not included in any cycle of $G$.
By (i), there are no two of these maximal cycles $C^{1},\cdots,C^{X}$ such that the alternatives included in them are identical. Moreover, by (ii), any alternative that is included in at least one cycle of $G$ is also included in some of these maximal cycles $C^{1},\cdots,C^{X}$.
We can construct $C^{1},\cdots,C^{X}$ satisfying (i) and (ii) by the following method. First, since we assume that there is one cycle, we can let $C^{1}$ be a maximal cycle of $G$. If $C^{1}$ satisfies (ii); that is, if $a\in A\setminus A^{1}$ is not included in any cycle of $G$, then $X=1;$ that is, only one maximal cycle is formed. If otherwise, then there must exists $a\in A\setminus A^{1}$ that is included in a cycle of $G$. Thus, in this case, there is a maximal cycle including $a\in A\setminus A^{1}$. Let $C^{2}$ be such a maximal cycle. Then, for $C^{1}$ and $C^{2},$ $A^{1}\setminus A^{2}\neq\emptyset$ and $A^{2}\setminus A^{1}\neq\emptyset,$ because $a\in A^{2}\setminus A^{1}$ and $C^{1}$ is a maximal cycle. Inductively, we can construct $C^{1},\cdots,C^{X}$ satisfying (i) and (ii).
Lemma 2
Suppose that there is a cycle of $G$. If Condition 1 is satisfied and $X\geq 2$, then, for all $x,y=1,\cdots,X$ such that $x\neq y$, $C^{x}\cap C^{y}=\emptyset$.
Proof. Toward a contradiction, suppose $ab\in C^{x}\cap C^{y}$. Since there is $v\in V$ such that $C^{x}\subseteq G_{v}$ by Condition 1, for all $c,d\in A^{x}$ such that $c\neq d$, $cd\in G$. Thus, there is a path from $a$ to $b$ with all alternatives in $A^{x}$. Similarly, there is a path from $b$ to $a$ with all alternatives in $A^{y}$. Therefore, there is a cycle including all alternatives in $A^{x}\cup A^{y}$. By (i), $A^{x}\subsetneq A^{x}\cup A^{y}$. Therefore, this fact contradicts that $C^{x}$ is a maximal cycle. Q.E.D.
Lemma 2 ensures that the ARF introduced below is well-defined.
By Condition 1, there is at least one individual $v$ such that $C^{x}\subseteq G_{v}$ for all $x=1,\cdots,X$. Thus, we let $v^{x}$ be such an individual for each $x=1,\cdots,X$. For each $x=1,\cdots,X$ and each $ab\in C^{x}$, let
$aP\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)b\text{ if }aP\left(R_{v^{x}}\right)b\text{ or [}aI\left(R_{v^{x}}\right)b\text{ and }aR^{T}b\text{]}$
$bP\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)a\text{ if otherwise.}$.
Note that if there is no cycle of $G$, then $A^{0}=A$ and thus $R^{\ast\ast}\left(\mathbf{R}\right)$ is empty in the construction so far.
Next, we consider each pair of alternatives $a$ and $b$ such that $a$ $\in A^{x}$ and $b\in A^{y}$ where $x,y=0,1,\cdots,X$ and either $x=y=0$ or $x\neq y$. If $V\left(a\right)\cap V\left(b\right)\neq\emptyset,$ then we choose $v\in V\left(a\right)\cap V\left(b\right)$ and let
$aP\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)b\text{ if }aP\left(R_{v}\right)b\text{ or [}aI\left(R_{v}\right)b\text{ and }aR^{T}b\text{]}$
$bP\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)a\text{ if otherwise.}$.
Otherwise; that is, if $V\left(a\right)\cap V\left(b\right)=\emptyset,$ then neither $aR^{\ast\ast}\left(\mathbf{R}\right)b$ nor $bR^{\ast\ast}\left(\mathbf{R}\right)a$.
By (ii), if $a\neq b$ and $V\left(a\right)\cap V\left(b\right)\neq\emptyset$, then either $aR^{\ast\ast}\left(\mathbf{R}\right)b$ or $bR^{\ast\ast}\left(\mathbf{R}\right)a$, and either of the two is determined by the evaluation of one individual (and the tie-breaking rule). Otherwise, then neither $aR^{\ast\ast}\left(\mathbf{R}\right)b$ nor $bR^{\ast\ast}\left(\mathbf{R}\right)a$. Therefore, $R^{\ast\ast}\left(\mathbf{R}\right)$ is asymmetric.
Finally, let
$f^{\ast\ast}\left(\mathbf{R}\right)$$\in$ $\mathcal{E}\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)\text{if }R^{\ast\ast}\left(\mathbf{R}\right)\text{ is acyclic}$
$=$ arbitrary weak order if otherwise.
Lemma 3
If Condition 1 is satisfied, then $f^{\ast\ast}$ is well-defined.
Proof. Suppose that Condition 1 is satisfied. By Lemma 2, for each $(a,b)$ satisfying $a\neq b$ and $V\left(a\right)\cap V\left(b\right)\neq\emptyset$, there is only one individual who determines either $aR^{\ast\ast}\left(\mathbf{R}\right)b$ or $bR^{\ast\ast}\left(\mathbf{R}\right)a$. Moreover, by Remark 1, if $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic, $\mathcal{E}\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)\neq\emptyset$. Thus, $f^{\ast\ast}$ is well-defined. Q.E.D.
We compare $f^{\ast\ast}$ with $f^{\ast}$, which is introduced earlier. We have $R^{\ast}\left(\mathbf{R}\right)\subseteq R^{\ast\ast}\left(\mathbf{R}\right)$ and thus $\mathcal{E}\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)\subseteq\mathcal{E}\left(R^{\ast}\left(\mathbf{R}\right)\right)$. Moreover, if $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic, then $R^{\ast}\left(\mathbf{R}\right)$ is also acyclic. Therefore, if $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic, then $f^{\ast\ast}\left(\mathbf{R}\right)\in\mathcal{E}\left(R^{\ast}\left(\mathbf{R}\right)\right)$; that is, $f^{\ast\ast}\left(\mathbf{R}\right)$ can be a result of $f^{\ast}\left(\mathbf{R}\right)$. This implies that we can regard $f^{\ast\ast}\left(\mathbf{R}\right)$ as $f^{\ast}\left(\mathbf{R}\right)$ with a complicated tie-breaking (incomparability-breaking) rule. We show that this rule ensures that the ARF satisfies IIA under Condition 1 later.
We provide an example to understand $f^{\ast\ast}$.
Example 1
There are seven alternatives $A=\left\{a_{1},\ldots,a_{7}\right\}$. There are only three types $1,$ $2$ and $3$ of individuals with regard to their evaluable set. Suppose that $v_{i}$ is a type $i=1,2,3$ individual. The evaluable sets of $1,2$ and $3$ are given by
$A_{v_{1}}$ $=$ $\{a_{1},a_{2},a_{3},a_{4}\},$
$A_{v_{2}}$ $=$ $\{a_{4},a_{5},a_{6}\},$
$A_{v_{3}}$ $=$ $\{a_{6},a_{7}\}.$
Moreover, the evaluable set of any other individual is either of the three. Then, $G$ is written in Figure 1.
Figure 1: $G$ of Example 1
Then, there are two maximal cycles
$C^{1}$ $=$ $\{a_{1}a_{2},a_{2}a_{3},a_{3}a_{4},a_{4}a_{1}\},$
$C^{2}$ $=$ $\{a_{4}a_{5},a_{5}a_{6},a_{6}a_{4}\},$
and $A^{0}=\left\{a_{7}\right\}$. We let $v^{1}=v_{1}$ and $v^{2}=v_{2};$ that is, relative positions of $a_{1},\cdots,a_{4}$ and those of $a_{4},a_{5},a_{6}$ are determined by $v_{1}$ and $v_{2},$ respectively. Moreover, we assume that relative positions of $a_{6}$ and $a_{7}$ is determined by $v_{3}$.
Now, let
$R_{v_{1}}$ ::: $a_{2},a_{4},a_{1},a_{3}$
$R_{v_{2}}$ ::: $a_{5},a_{6},a_{4},$
$R_{v_{3}}$ ::: $a_{7},a_{6}$
Then,
$R^{\ast\ast}\left(\mathbf{R}\right)$ $=$ $R_{v_{1}}\cup R_{v_{2}}\cup R_{v_{3}}$
$=$ $\{a_{2}a_{4},a_{2}a_{1},a_{2}a_{3},a_{4}a_{1},a_{4}a_{3},$
$a_{1}a_{3},a_{5}a_{6},a_{5}a_{4},a_{6}a_{4},a_{7}a_{6}\},$
which is acyclic. Then, since the outcome of this function is a linear order extension of $R^{\ast\ast}\left(\mathbf{R}\right)$, for example,
$f^{\ast\ast}\left(\mathbf{R}\right):a_{2},a_{5},a_{7},a_{6},a_{4},a_{1},a_{3}\text{.}$
We introduce the following result.
Theorem 2
There is an ARF that satisfies TV, PC, and IIA if and only if Condition 1 is satisfied.
Proof. By Theorem 1, Condition 1 is trivially necessary. Thus, we show the if-part only. We show that if Condition 1 is satisfied, then $f^{\ast\ast}$ satisfies TV, PC, and IIA.
First, we show the following result.
Claim 2
If Condition 1 is satisfied, then $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic for any $\mathbf{R}\in\mathcal{W}$.
Proof. Suppose not; that is, there are distinct $a_{1},a_{2},\cdots,a_{M}\in A$ such that $a_{m+1}R^{\ast\ast}\left(\mathbf{R}\right)a_{m}$ for all $m=1,\cdots,M$. Then, there exists a cycle that is defined by (1) and thus there also exists $C^{x}$ such that $\left\{a_{1},a_{2},\cdots,a_{M}\right\}\subseteq A^{x}$. Then, for all $m=1,\cdots,M$, either $a_{m+1}P\left(R_{v^{x}}\right)a_{m}$ or [$a_{m+1}I\left(R_{v^{x}}\right)a_{m}$ and $a_{m+1}R^{T}a_{m}$]. Therefore, in any case, $a_{m+1}R_{v^{x}}a_{m}$ for all $m=1,\cdots,M$. By the transitivity of $R_{v^{x}}$ and $a_{M+1}=a_{1}$, $a_{m}R_{v^{x}}a_{m+1}$ for all $m=1,\cdots,M$. Hence, $a_{m+1}I\left(R_{v^{x}}\right)a_{m}$ and thus $a_{m+1}R^{T}a_{m}$ for all $m=1,\cdots,M$, but it contradicts the transitivity of $R^{T}$. Therefore, $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic. Q.E.D.
Second, we show that if $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic, $f^{\ast\ast}$ satisfies IIA. Let $a,b\in A$ such that $a\neq b$ and $V\left(a\right)\cap V\left(b\right)\neq\emptyset$, $\mathbf{R}$ and $\mathbf{R}^{\prime}$ such that $\left.R_{v}\right|_{\{a,b\}}=\left.R_{v}^{\prime}\right|_{\{a,b\}}$ for all $v\in V\left(a\right)\cap V\left(b\right)$. Then, by the construction, there is $v^{\ast}\in V\left(a\right)\cap V\left(b\right)$ such that
$aP\left(f^{\ast\ast}\left(\mathbf{R}\right)\right)b\text{ if }aP\left(R_{v^{\ast}}\right)b\text{ or [}aI\left(R_{v^{\ast}}\right)b\text{ and }aR^{T}b\text{]}$
$bP\left(f^{\ast\ast}\left(\mathbf{R}\right)\right)a\text{ if otherwise.}$.
Since $\left.R_{v^{\ast}}\right|_{\{a,b\}}=\left.R_{v^{\ast}}^{\prime}\right|_{\{a,b\}}$, $\left.f^{\ast\ast}\left(\mathbf{R}\right)\right|_{\{a,b\}}=\left.f^{\ast\ast}\left(\mathbf{R}^{\prime}\right)\right|_{\{a,b\}}$. Therefore, $f^{\ast\ast}$ satisfies IIA.
Finally, we show that if $R^{\ast\ast}\left(\mathbf{R}\right)$ is acyclic, then $f^{\ast\ast}$ satisfies PC. If $aP\left(R_{v}\right)b$ for all $v\in V\left(a\right)\cap V\left(b\right)$, then $aP\left(R^{\ast\ast}\left(\mathbf{R}\right)\right)b$. Therefore, $f^{\ast\ast}$ satisfies PC. Then, by Claim 2, if Condition 1 is satisfied, then $f^{\ast\ast}$ satisfies TV, PC, and IIA. Q.E.D.
Condition 1 guarantees the existence of an ARF that satisfies the three properties. Thus, the if-part of Theorem 1 is generalized. On the other hand, the only-if-part of Theorem 2 is implied by that of Theorem 1.
We consider some sufficient conditions for the existence of an ARF $f$ that satisfies the three properties in order to clarify the result.
Corollary 1
If $G$ has no cycle, then there is an ARF that satisfies TV, PC, and IIA.
This result implies that if $\mathbf{A}$ is sufficiently small, then there exists an ARF that satisfies the three properties. We briefly show this fact. Let $\mathbf{A}^{\prime}$ be a smaller evaluability profile than $\mathbf{A}$ and $G^{\prime}$ be the graph constructed from the problem $(V,A,\mathbf{A}^{\prime})$. Then, $G^{\prime}\subseteq G$. Thus, if $G$ has no cycle, then $G^{\prime}$ also has no cycle.
Corollary 2
If there is at least one complete individual, then there is an ARF that satisfies TV, PC, and IIA.
This result is trivial from Lemma 1 and Theorem 2. This implies that if an evaluability profile $\mathbf{A}$ is sufficiently large, then there is an ARF $f$ that satisfies the three properties. This is because, if there is a complete individual, then the individual must also be complete for a larger evaluability profile.
By Corollaries 1 and 2, if $\mathbf{A}$ is either sufficiently small or sufficiently large, then there is an ARF that satisfies the three properties. Conversely, if $\mathbf{A}$ is neither very small nor very large, then an ARF $f$ that satisfies TV and PC may fail to exist.
Next, we focus on the situation where Condition 2 is not satisfied.
Corollary 3
Suppose that Condition 2 is not satisfied. Then, there is an ARF that satisfies TV, PC, and IIA if and only if there is at least one complete individual. Moreover, this is satisfied even without IIA.
This result is straightforward from Lemma 1, Theorems 1 and 2. Note that in the peer rating model, there is a Hamiltonian cycle and there is no complete individual. Thus, as also shown by Ando et al. (2003), there is no ARF that satisfies TV and PC in the model.
Third, we additionally require ND. We first consider the case where at least one individual is complete. By Corollary 2, in this case, there is an ARF that satisfies TV, PC, and IIA. Ohbo et al. (2005) show the following result.
Remark 3
Suppose that there is at least one complete individual. If an ARF $f$ satisfies TV, PC, and IIA, then there is a dictator of $f$.
Since all individuals are complete in the model of Arrow (1963), Remark 3 is a generalization of Arrow’s impossibility result. Here, we further generalize Remark 3.
Theorem 3
There is an ARF that satisfies TV, PC, IIA, and ND if and only if Conditions 1 and 2 are satisfied.
Proof. We show the only-if-part. By Remark 4, it is necessary that no complete individual exists. By Corollary 3, if there is no complete individual but $G$ has a Hamiltonian cycle, then there is no ARF $f$ satisfies TV and PC. Thus, Condition 2 is a necessary condition. Moreover, by Theorem 2, Condition 1 is also necessary.
Next, show the if-part. Since there is no Hamiltonian cycle, any individual $v$ is not complete. Hence $f^{\ast\ast}$ must satisfies ND. Thus, by this fact and Theorem 2, we have the if-part of Theorem 3. Q.E.D.
By this result, if we require ND in addition to others, then Condition 2 is a necessary (and sufficient) additional condition to ensure the existence of an ARF satisfying all of them. By Corollary 2, there is an ARF satisfying them excepting ND if $\mathbf{A}$ is sufficiently large. However, if ND is additionally required, then there is no ARF satisfying them if $\mathbf{A}$ is sufficiently large. In other words, when an evaluability profile $\mathbf{A}$ is large as to have a Hamiltonian cycle, an ARF that satisfies TV, PC, and IIA, must be a dictatorship rule.
On the other hand, we immediately have the following result.
Corollary 4
If $G$ has no cycle, then there is an ARF that satisfies TV, PC, IIA, and ND.
This result is stronger than Corollary 1 and implies that if $\mathbf{A}$ is sufficiently small, then there is an ARF other than the dictatorship ones that satisfies TV, PC, and IIA.
Next, we focus on the class of problems where the evaluability profiles are non-trivial. This class is discussed by Fujiwara-Greve et al. (2023). In this class, we have the following impossibility result.
Corollary 5
If $A$ is non-trivial, then there is no ARF that satisfies TV, PC, IIA, and ND.
As noted above, $A$ is non-trivial if and only if $G$ is complete. First, if there is no complete individual, then Condition 1 is not satisfied. By Theorem 2, in this case, there is no ARF that satisfies TV, PC, and IIA. Second, if there is a complete individual, then, by Remark 4, any ARF $f$ satisfying TV, PC, and IIA does not satisfy ND. Therefore, we have Corollary 5.
Next, we consider a wPC and NC instead of PC, because they are weaker than PC because of Remark 1. First, we have the following result.
Lemma 4
Suppose that an ARF $f$ satisfies TV, wPC, and IIA, and Condition 1 is not satisfied; that is, there is a cycle of $G$ that is defined by (1) without $v\in V$ such that $C\subseteq G_{v}$. For any $\mathbf{R}\in\mathcal{W},$ and any $l,n\in\left\{1,\cdots,M\right\}$ such that $l\neq n$, $a_{l}I\left(f\left(\mathbf{R}\right)\right)a_{n}$.
Proof. This proof is inspired by that of Ando et al. (2003, Theorem 3.7). First, we show the following result.
Claim 3
Suppose that an ARF $f$ satisfies TV, wPC, and IIA, and Condition 1 is not satisfied; that is, there is a cycle of $G$ that is defined by (1) without $v\in V$ such that $C\subseteq G_{v}$. For any $n=1,\cdots,M$, if $a_{n+1}P\left(R_{v}\right)a_{n}$ for all $v\in V\left(a_{n}\right)\cap V\left(a_{n+1}\right)$, then $a_{n+1}I\left(f\left(\mathbf{R}\right)\right)a_{n}$, where $a_{M+1}=a_{1}$.
Proof. We can let $\mathbf{\bar{R}=}\left(\bar{R}_{v}^{C}\right)_{v\in V}$ be a cyclic profile of $C$ that is defined in the proof of Theorem 1, because there is no $v\in V$ such that $C\subseteq G_{v}$. Then, for all $m=1,\cdots,M$, $a_{m+1}P\left(\bar{R}_{v}\right)a_{m}$ for all $v\in V\left(a_{m}\right)\cap V\left(a_{m+1}\right)$. Since $f$ satisfies wPC, for any $n=1,\cdots,M,$ $a_{n+1}f\left(\mathbf{\bar{R}}\right)a_{n}$. Then, by the transitivity of $f\left(\mathbf{\bar{R}}\right)$ and $a_{M+1}=a_{1}$, we also have $a_{n}f\left(\mathbf{\bar{R}}\right)a_{n+1}$ and thus $a_{n+1}I\left(f\left(\mathbf{\bar{R}}\right)\right)a_{n}$ for any $n=1,\cdots,M$.
Now, we consider an arbitrary $\mathbf{R}$ such that $a_{n+1}P\left(R_{v}\right)a_{n}$ for all $v\in V\left(a_{n}\right)\cap V\left(a_{n+1}\right)$. Then, since
$\left.R_{v}\right|_{\{a_{n},a_{n+1}\}}=\left.\bar{R}_{v}\right|_{\{a_{n},a_{n+1}\}}$
for all $v\in V$ and $f$ satisfies IIA, we have $a_{n+1}I\left(R_{v}\right)a_{n}$. Q.E.D.
Next, we show that for any $\mathbf{R}\in W_{V},$ and for any $n=1,\cdots,M$, $a_{n+1}I\left(f\left(\mathbf{R}\right)\right)a_{n}$, where $a_{M+1}=a_{1}$ and $A^{C}=\left\{a_{1},\cdots,a_{M}\right\}$ be the set of alternatives that are included in cycle $C$ defined in Lemma 4.
Without loss of generality, we let $n=1$ and show $a_{2}I\left(f\left(\mathbf{R}\right)\right)a_{1}$. Let $\mathbf{R}$ be an arbitrary preference profile of individuals. We introduce another preference profile of individuals denoted by $\mathbf{R}^{\prime}$ that have the following characteristics. First, let $R^{\ast}$ be a linear order on $A^{C}$ satisfying
$R^{\ast}:a_{M},a_{M-1},\cdots,a_{2},a_{1}.$
Then, let $\mathbf{R}^{\prime}$ be such that (i) $\left.R_{v}^{\prime}\right|_{\{a_{1},a_{2}\}}=\left.R_{v}\right|_{\{a_{1},a_{2}\}}$ for all $v\in V\left(a_{n}\right)\cap V\left(a_{n+1}\right)$, (ii) for all $m\in\{2,\cdots,M\}$
$\left.R_{v}^{\prime}\right|_{\{a_{m},a_{m+1}\}}=\left.R^{\ast}\right|_{\{a_{m},a_{m+1}\}}$
for all $v\in V\left(a_{m}\right)\cap V\left(a_{m+1}\right)$. Since $a_{l}P\left(R^{\ast}\right)a_{2}$ and $a_{l}P\left(R^{\ast}\right)a_{1}$ for all $l\in\{3,\cdots,M\}$, $a_{2}R_{v}^{\prime}a_{1}$ or $a_{1}R_{v}^{\prime}a_{2}\,$(or both) is possible. Therefore, $\mathbf{R}^{\prime}$ is well-defined.
By Claim 3, $a_{m+1}I\left(f\left(\mathbf{R}^{\prime}\right)\right)a_{m}$ for all $m\in\{2,\cdots,M\}$. Therefore, by the transitivity of $I\left(f\left(\mathbf{R}^{\prime}\right)\right)$, $a_{2}I\left(f\left(\mathbf{R}^{\prime}\right)\right)a_{1}$. Moreover, since $\left.R_{v}^{\prime}\right|_{\{a_{1},a_{2}\}}=\left.R_{v}\right|_{\{a_{1},a_{2}\}}$ for all $v\in V\left(a_{m}\right)\cap V\left(a_{m+1}\right)$ and $f$ satisfies IIA, we have $a_{2}I\left(f\left(\mathbf{R}\right)\right)a_{1}$.
Then, for any $\mathbf{R}\in W_{V},$ and for any $n=1,\cdots,M$, $a_{n+1}I\left(f\left(\mathbf{R}\right)\right)a_{n}$, where $a_{M+1}=a_{1}$. By the transitivity of $I\left(f\left(\mathbf{R}\right)\right)$, for any $l,n\in\left\{1,\cdots,M\right\}$ such that $l\neq n$, $a_{l}I\left(f\left(\mathbf{R}\right)\right)a_{n}$. Q.E.D.
By Lemma 4, we have the following result.
Theorem 4
There is an ARF $f$ satisfies TV, wPC, NC, and IIA if and only if Condition 1 is satisfied.
Proof. By Theorem 2, the only-if-part is trivial.
We show the if-part. By Lemma 4, if Condition 1 is not satisfied; that is, if there is a cycle of $G$ that is defined by (1) where there is no $v\in V$ such that $C\subseteq G_{v}$, then any two alternatives included in the cycle are always indifferent with any ARF satisfying TV, wPC, and IIA. Therefore, in this case, any ARF satisfying TV, wPC, and IIA does not satisfy NC. Q.E.D.
Thus, even if we require wPC and NC instead of PC, Condition 1 is still necessary. Ando et al. (2003) show that in the peer rating model, there is no ARF $f$ that satisfies TV, NC, and IIA. The necessary and sufficient condition that there is an ARF $f$ that satisfies TV, NC, and IIA is unknown. Ohseto (2007) also examines the peer rating model and shows that the Borda rule satisfies TV, wPC, and NC, but it does not satisfy IIA. [^8] Ohseto (2007) does not mention that the rule satisfies NC, but it is trivially satisifed.
4 Proportions of evaluability profiles
By Theorems 2 and 3, we can classify the evaluability profile into three types. First one is that for which there is no ARF that satisfies TV, PA, and IIA; that is, neither Condition 1 nor 2 is satisfied with this type of profile. Hereafter, we refer this type of evaluability profile as an “impossible” profile (hereafter IP). Second one is that for which an ARF that satisfies TV, PA, and IIA must be a dictatorship rule; that is, Condition 1 is satisfied but Condition 2 is not with this type of profile. Hereafter, we refer this type of evaluability profile as a dictatorship profile (hereafter DP). Third one is that for which there is an ARF that satisfies TV, PA, IIA, and ND; that is, Conditions 1 and 2 are satisfied with this type of profile. To discuss how frequently each type occurs, we derive the proportion of each type. Hereafter, we refer this type of evaluability profile as a “possible” profile (hereafter PP). Note that an IP is such that TV, PA, and IIA are impossible, and a DP is such that ND as well as the three are possible.
We derive all possible evaluable sets of an individual. For example, suppose $A=\left\{a,b,c\right\}$. Then, the potential evaluable sets for an individual are $\left\{a,b,c\right\}$, $\left\{a,b\right\}$, $\left\{a,c\right\}$, and $\left\{b,c\right\}$, because each of them must include at least two alternatives. In general, there are $12^{\left|A\right|}-\left|A\right|-$ 1 potential evaluable sets for each individual. Therefore, in total, there are $\left(2^{\left|A\right|}-\left|A\right|-1\right)^{\left|I\right|}$ potential evaluability profiles.
First, we consider the proportion of the DPs. By Lemma 1, in this type, there is at least one complete individual for the evaluability profile. Therefore, there are DPs that account for
$1-\left(\frac{2^{\left|A\right|}-\left|A\right|-2}{2^{\left|A\right|}-\left|A\right|-1}\right)^{\left|I\right|}$ (2)
of the total, and the remainder consists of either IPs or PPs. In each cell of the following table, the approximation of (2) with the corresponding combination of $\left|I\right|$ and $\left|A\right|$ is written. By the table, if $\left|A\right|$ is small, then in many evaluability profiles, there is at least one complete individual and the ARF satisfying TV, PA, and IIA must be a dictatorship rule. On the other hand, if $\left|A\right|$ is large, then almost evaluability profiles is such that there is no ARF satisfying TV, PA, and IIA or there is some ARF satisfying TV, PAIIA, and ND.
$\left|A\right|\setminus\left|I\right|$ 3 6 9 12 15 18 21 24 27
3 0.58 0.82 0.92 0.97 0.99 0.99 1 1 1
5 0.11 0.21 0.30 0.38 0.44 0.51 0.56 0.61 0.65
7 0.02 0.05 0.07 0.10 0.12 0.14 0.16 0.18 0.20
9 0.00 0.01 0.02 0.02 0.03 0.04 0.04 0.05 0.05
Table 1: Proportions of DPs
Deriving the proportion of IPs and that of PPs is relatively difficult, even if neither $\left|A\right|$ nor $\left|I\right|$ is very large. Here, we derive them only the three cases where $\left(\left|I\right|,\left|A\right|\right)=\left(3,3\right),$ $\left(4,3\right)$ and $\left(3,4\right)$.
First, if $\left(\left|I\right|,\left|A\right|\right)=\left(3,3\right),$ then there are $6$, $37$, and $21$ evaluability profiles that are IPs, DPs, and PPs, respectively. Thus, about $33\%$ of all profiles are for which there is an ARF that satisfies TV, PAIIA, and ND. Second, if $\left(\left|I\right|,\left|A\right|\right)=\left(4,3\right),$ then there are $36$, $175$, and $45$ evaluability profiles that are IPs, DPs, and PPs, respectively. Thus, about $18\%$ of all profiles are for which there is an ARF that satisfies the four properties. By comparing the two cases, the proportion of PPs may decrease with the number of individuals. Third, if $\left(\left|I\right|,\left|A\right|\right)=\left(3,4\right),$ then there are $371$, $331$, and $629$ evaluability profiles that are IPs, DPs, and PPs, respectively. Thus, about $47\%$ of all profiles are for which there is an ARF that satisfies the four properties. By comparing the cases where $\left(\left|I\right|,\left|A\right|\right)=\left(3,3\right)$ and $\left(3,4\right)$, the proportion of PPs may increase with the number of individuals.
While this analysis provides only limited conclusions, it suggests that when the number of individuals is much larger than that of alternatives, requiring TV, PA, IIA, and ND may be almost impossible, whereas when the number of alternatives is much larger than that of individuals, it may be almost possible.
5 Concluding Remarks
In the social welfare function à la Arrow, it is assumed that each individual evaluates all alternatives. Moreover, in real voting systems such as national elections, we may be supposed to vote after evaluating all alternatives. However, we are sometimes not well acquainted enough to evaluate all alternatives.
Here, we consider a situation where some individuals lack information about a certain alternative, while other individuals are well-informed about and prefer that alternative. Moreover, the other alternatives are known by all individuals. In such a case, since individuals are likely to be risk-averse, many of those who are not well-informed about the alternative may express their preferences by ranking it lower than other alternatives. Consequently, that alternative would end up with a low position in the aggregated ranking. On the other hand, as discussed in this study, if no individual evaluates their unknown alternatives, then the alternative may be ranked higher in the aggregated ranking.
Therefore, although whether or not the rule “no individual evaluates their unknown alternatives” should be adopted is debatable, the decision to adopt this rule may have a significant impact on the aggregated ranking.
Reference
Adachi, T., 2014. A natural mechanism for eliciting ranking when jurors have favorites. Games and Economic Behavior 87, 508–518.
Amorós, P., 2009. Eliciting socially optimal rankings from unfair jurors. Journal of Economic Theory 144, 1211–1226.
Amorós, P., 2011. A natural mechanism to choose the deserving winner when the jury is made up of all contestants. Economics Letters 110, 241–244.
Amorós, P., 2019. Choosing the winner of a competition using natural mechanisms: Conditions based on the jury, Mathematical Social Sciences 98 26–38
Amorós, P., Corchón, L., Moreno, B., 2002. The scholarship assignment problem. Games and Economic Behavior 38, 1–18.
Ando, K., Ohara, A., Yamamoto, Y. 2003. Impossibility theorems on mutual evaluation (in Japanese), Journal of the Operations Research Society of Japan 46, 523–532.
Ando, K., Tsurutani, M. Umezawa, M., Yamamoto, Y. 2007. Impossibility and possibility theorems for social choice functions on incomplete preference profiles, Pacific Journal of Optimization 3(1), 11-25.
Ariely, D. 2010. The Upside of Irrationality: The Unexpected Benefits of Defying Logic at Work and at Home, HarperCollins, New York.
Arrow, K.J., 1963. Social Choice and Individual Values, 2nd Edition. Wiley, New York.
Barberá, S., Bossert W. 2023. Opinion aggregation: Borda and Condorcet revisited, Journal of Economic Theory 210, 105654.
Cengelci, M. A., Sanver, M. R. 2007. Is abstention an escape from Arrow’s theorem? Social Choice and Welfare 28, 439–442.
Duggan, J. 1999. A General Extension Theorem for Binary Relations, Journal of Economic Theory 86, 1-16.
Edelman, P.H., Por, A., 2021. A new axiomatic approach to the impartial nomination problem. Games and Economic Behavior 130, 443–451.
Fujiwara-Greve, T., Kawada, Y., Nakamura, Y., Okamoto, N., 2023. Accountable Voting. Available at SSRN: https://ssrn.com/abstract=4406802
Hansson, B. 1973. The independence condition in the theory of social choice. Theory and Decision 4, 25–49.
Holzman, R., Moulin, H., 2013. Impartial nominations for a prize. Econometrica 81, 173-196.
Kitahara, M., Okumura, Y. 2023. On Extensions of Partial Priorities in School Choice. Mimeo Available at SSRN: https://ssrn.com/abstract=4462665
Mackenzie, A., 2015. Symmetry and impartial lotteries. Games and Economic Behavior 94, 15–28.
Mackenzie, A., 2019. Impartial nominations for a prize. Economic Theory 69, 713–743.
Ng, Y.K., Sun, G.Z. 2003. Exclusion of self evaluations in peer ratings: an impossibility and some proposals, Social Choice and Welfare 20, 443–456.
Ohseto, S. 2007. A characterization of the Borda rule in peer ratings, Mathematical Social Sciences 54(2) 147-151.
Ohbo, K., Tsurutani, M., Umezawa, M., Yamamoto, Y. 2005. Social welfare function for restricted individual preference domain, Pacific Journal of Optimization 1(2), 315-325.
Quesada, A. 2005. Abstention as an escape from Arrow’s theorem. Social Choice and Welfare, 25 221–226.
Tamura, S. 2016. Characterizing minimal impartial rules for awarding prizes, Games and Economic Behavior 95, 41-46.
Tamura, S., Ohseto, S., 2014. Impartial nomination correspondences, Social Choice and Welfare 43, 47-54.
Villar, A., 2023. The precedence function: a numerical evaluation method for multi criteria ranking problems, Economic Theory Bulletin 11, 211–219.